OpenStudy (anonymous):

WILL FAN AND GIVE MEDAL AND TESTIMONY!!! The one-sample t statistic for a test of Ho: µ = 10 Ha: µ <10 Based on n = 10 observations has the value t = -2.25. a. What are the degrees of freedom for this statistic? b. What is the P-value for this test?

OpenStudy (anonymous):

Degrees of freedom: n-1=10-1=9